OpenStudy (anonymous):

m^2-1/2m=7/2....solve by completing the square

OpenStudy (amistre64):

id multiply by 2 to get rid of fractions then divide it out in the end

OpenStudy (amistre64):

m^2-1/2m=7/2 2m^2-1m=7 2(m^2-1/2m)=7 2(m^2-1/2m+(1/2)^2)=7 + 2(1/2)^2 and so forth

OpenStudy (amistre64):

i spose it makes no difference but its just easier for me to work with for some reason